Week of 171009

WLCG Operations Call details

  • At CERN the meeting room is 513 R-068.

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Web
  • Whenever a particular topic needs to be discussed at the daily meeting requiring information from site or experiments, it is highly recommended to announce it by email to wlcg-operations@cernSPAMNOTNOSPAMPLEASE.ch to make sure that the relevant parties have the time to collect the required information or invite the right people at the meeting.

Tier-1 downtimes

Experiments may experience problems if two or more of their Tier-1 sites are inaccessible at the same time. Therefore Tier-1 sites should do their best to avoid scheduling a downtime classified as "outage" in a time slot overlapping with an "outage" downtime already declared by another Tier-1 site supporting the same VO(s). The following procedure is recommended:

  1. A Tier-1 should check the downtimes calendar to see if another Tier-1 has already an "outage" downtime in the desired time slot.
  2. If there is a conflict, another time slot should be chosen.
  3. In case stronger constraints cannot allow to choose another time slot, the Tier-1 will point out the existence of the conflict to the SCOD mailing list and at the next WLCG operations call, to discuss it with the representatives of the experiments involved and the other Tier-1.

As an additional precaution, the SCOD will check the downtimes calendar for Tier-1 "outage" downtime conflicts at least once during his/her shift, for the current and the following two weeks; in case a conflict is found, it will be discussed at the next operations call, or offline if at least one relevant experiment or site contact is absent.

Links to Tier-1 downtimes

ALICE ATLAS CMS LHCB
  BNL FNAL  

Monday

Attendance:

  • local: Kate (DB, WLCG, chair), Julia (WLCG), Maarten (WLCG, ALICE), Gavin (computing), Belinda (storage), Alberto (monitoring), Mayank (WLCG), Tomas J. (ATLAS), Marian (network)
  • remote: Elena (CNAF), Herve (storage), Ulf (NDGF), Chi (ASGC), Christoph (CMS), David (IN2P3), Gareth (RAL), Dave M (FNAL), Pepe (PIC), Kyle (OSG), Sang-Un (KISTI), Xin (BNL), Vincenzo (EGI), Andrew (NL-T1), Victor (JINR)

Experiments round table:

  • ATLAS reports ( raw view) -
    • Stabily around 250k of running slots of jobs
    • EOS
      • 1 file lost on CERN-PROD_DATADISK
        • Still needs investigation
        • Recreated from other replicas
      • Traffic to EOS was suffering from timeouts when calculating checksums.
        • fixed quickly by eos experts on Friday
    • NDGF-T1:
      • SCRATCHDISK got full, quota increased
      • TAPE buffers were close to full on Sunday

  • CMS reports ( raw view) -
    • High cpu utilization: ~180k core (production) + ~35k cores (analysis) for most over last week
    • Transfer system also under high pressure
      • Presently over 1PB backlog of data to be transferred from CERN to T1s
    • KIT Disk almost full: GGUS:130929
      • Lowered threshold for DDM to use less disk
      • Provided a few orphan files to clean
    • Castor at RAL went out of disk space: GGUS:130949 and GGUS:130941
      • Lowered thresholds for DDM to use less disk
      • Consistency check to identify possible orphan files in preparation
Gareth commented that the diskservers were drained but are now back.

  • ALICE -
    • Low to high activity on average
      • Low activity due to CERN resources draining for kernel security patch
    • CERN: many expired jobs due to 1 HTCondor CE going AWOL
    • CERN: EOS-ALICE ~30 min downtime Thu afternoon (OTG:0040164)
    • CERN: many reco job failures on the Oracle cloud resources
      • Due to absence of HEP_OSlibs_SL6 on the WN (fixed)
Christoph asked if RPM was missing from the beginning or was removed. Maarten confirmed it was never there. Gavin mentioned that the Docker image had to be updated.

  • LHCb reports ( raw view) -
    • Activity
      • Monte Carlo simulation, data processing and user analysis
      • pre-staging of 2015 data for reprocessing is started and will continue during weeks.
    • Site Issues
      • T1:
        • Failures in transfers to and from RRCKI over the weekend, solved now.
        • NL-T1 worker nodes in downtime tomorrow and Wednesday.
        • Failed transfers from IC to SARA (IPV6) (GGUS:129946); Problem probably in SARA connection to LHCOne

Sites / Services round table:

  • ASGC: ntr
  • BNL: BNL is becoming a T1 site for Belle experiment.
  • CNAF:
    • ATLAS is to the pledge
    • LHCb will be to the pledge as soon as
    • For ATLAS and LHCb there will be a difference between accounting visible and real due to mirroring
    • We have problems with transfers tape for CMS
  • EGI: ntr
  • FNAL: ntr
  • IN2P3: ntr
  • JINR: A few errors in SAM at Buffer/MSS SE after lost of power on Ethernet switch 03-10-2017.
  • KISTI: ntr
  • KIT: nc
  • NDGF:We have a warning downtime for tomorrow due to rolling kernel updates (centos6)
    • atlas tape almost filled buffers during the weekend due to TSM software updates that required config changes
  • NL-T1: A network upgrade planned for tomorrow evening. 20 minute of network interruptions to be expected.
    • SARA - WNs will be upgraded tomorrow and the day after tomorrow, downtime was announced
  • NRC-KI: NC
  • OSG: nc
  • PIC: We had an incident in our Tape Robot last Friday. We stopped the RW access, meanwhile the problem was fixed. It was a hardware problem (broken elevator). It was promptly fixed, and then the RW access was granted again. Files sent to PIC were sent to Buffer and then later copied. The READS were all affected, but the problem was fixed in a few hours. Retries resumed well, so we hope it was not impacting badly to the VOs (none complained, afaik).
  • RAL: Disk servers were being drained from CMSDisk - in anticipation of move to Echo (CEPH). This led to the CMS disk space problem. Some disk servers put back into service on Friday to give CMS more space.
  • TRIUMF: NTR

  • CERN computing services: Second half of resources is being drained for OS upgrade.
  • CERN storage services:
    • Tuned a sub-optimal GridFTP timeout (Control-Connection Timeout) after some failed FTS transfers for Atlas end of last week. It turns out that for big files and/or slow remotes, the transfer might take more that the default 10 minutes. The transfer will succeed but when FTS wants to validate the checksum of the written file, the control connection has timed out, which marks the transfer as failed.
  • CERN databases: CMSONR ADGs needed to be restarted on Monday. Bug investigation is ongoing.
  • GGUS: ntr
  • Monitoring:
    • Draft reports for 09/2017 availability distributed by WLCG Office last Friday (6 October). 10 days for comments.
  • Networks: SARA issue is now solved
  • Security: nc

AOB:

Edit | Attach | Watch | Print version | History: r15 < r14 < r13 < r12 < r11 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r15 - 2017-10-09 - MaartenLitmaath
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback